OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 10x – 4 = 2y 2y – 10x = 2

OpenStudy (anonymous):

10x – 4 = 2y ----- (I) rearrange it to get 2y + 4 x = --------- 10 Now substitute it in (II) 2y – 10x = 2 ----- (II) and solve. You will get value of y. Substitute that value of y in (I) and get value of x.

OpenStudy (anonymous):

thanks...☺